Q.4) If an object starts out at rest and falls a distance of s feet in t seconds, then physical laws imply that s = 16t². Find ∆s/∆t as t changes from t0 to t0 + ∆t.



Use the result to find ∆s/∆t as t changes from:



(a) 3 to 3.5, (b) 3 to 3.2, and (c) 3 to 3.1

1
Expert's answer
2022-04-27T17:08:36-0400
"\Delta s=s(t_0+\Delta t)-s(t_0)"

"=16(t_0+\Delta t)^2-16(t_0)^2"

"=32t_0\Delta t+16(\Delta t)^2"

"\Delta s/\Delta t=32t_0+16\Delta t"

(a) "t_0=3 , \Delta t=3.5-3=0.5"


"\Delta s/\Delta t=32(3)+16(0.5)=104"

(b) "t_0=3 , \Delta t=3.2-3=0.2"


"\Delta s/\Delta t=32(3)+16(0.2)=99.2"

(c) "t_0=3 , \Delta t=3.1-3=0.1"


"\Delta s/\Delta t=32(3)+16(0.1)=97.6"